please show all of your work for parts b and c! thank you :)

  1. Your boss is studying a pepsin homolog present in fruit and asks you to make 1.00 L of a 0.100 M acetate buffer at pH 4.45. You have powdered sodium acetate and glacial (melting point 17 °C) acetic acid (also distilled water and all the usual lab equipment) available. When finalizing your answer, think about how best to measure the various components of the buffer. Note that the pKa of acetic acid is 4.79.

  1. What do you mix in what quantity to make the buffer? In answering this question, think about the most practical way to measure each of the materials you need to make the indicated buffer. (5 pts)

  1. Upon adding 0.010 millimoles of a strong acid to a 1.0 ml sample of your buffer, what is the new pH? (3 pts)
  1. For comparison, what is the pH of a 0.010 M aqueous solution of acetic acid? Of 0.010 M perchloric acid (a strong acid)? (3 pts)
Answer #1

a) First we find 0.1 m of 1 L buffer preparation

The density of water = 1g/ml

So 0.1 m = 0.1 M

Now as we know that

pH = pKa + log([base]/[acid])

4.45 = 4.79 + log([acetate]/[acetic acid])

[aceate] =0.457[acetic acid]

[acetic acid] + [acetate] = 0.1 M x 1 L = 0.1 mol

[acetic acid] + 0.45[acetic acid] = 0.1

Therefore the amount of

[acetic acid] = 0.066 mols

also,

[acetate] = 0.1 - 0.069

= 0.031 m


b)The moles of HCl added = 0.02 mmol

So the formed acetic acid = 0.1 x 1 + 0.01/1 = 0.11 M

= 0.1 M x 1 ml - 0.01 mmol/1 = 0.09 M

new pH = 4.79 + log(0.09/0.11) = 4.70

c) pH of 0.01 M acetic acid

Ka = 1.62 x 10^-5 = x^2/0.01

x = [H+] = 4.02 x 10^-4 M

pH = 3.39

pH of 0.01 M HCl

[H+] = 0.01 M

pH = 2